# Amuro & Char: Analyzing the Relationship between Pre-Training and Fine-Tuning of Large Language Models

### v2

@@ Abstract
The development of large language models leads to the formation of a pre-train-then-align paradigm, in which the model is typically pre-trained on a large text corpus and undergoes a tuning stage to align the model with human preference or downstream tasks. In this work, we investigate the relationship between pre-training and fine-tuning by fine-tuning multiple intermediate pre-trained model checkpoints. Our results on 18 datasets suggest that i) continual pre-training improves the model in a latent way that unveils after fine-tuning; ii) with extra fine-tuning, the datasets that the model does not demonstrate capability gain much more than those that the model performs well during the pre-training stage; iii) although model benefits significantly through supervised fine-tuning, it may forget previously known domain knowledge and the tasks that are not seen during fine-tuning; iv) the model resembles high sensitivity to evaluation prompts after supervised fine-tuning, but this sensitivity can be alleviated by more pre-training.  [^1] Code, results, and data to reproduce the experiments are available at https://anonymous.4open.science
/r/AmuroCharRelease-DEC5
. All the model checkpoints resulting from this work are available at https://huggingface.co/KaiserWhoLearns/PTvsSFT_OLMo1b
@@ 1 Introduction
The rise of large language models (LLMs) as a general-purpose tool for a diverse range of natural language processing tasks has dramatically transformed the field, introducing new paradigms for data collection and model training (Brown et al., 2020, Biderman et al., 2023, Touvron et al., 2023, Jiang et al., 2023, Chowdhery et al., 2023, Groeneveld et al., 2024, Wang et al., 2024, inter alia). Numerous models, training methods, datasets, and evaluation methods continue to be developed on an ongoing basis. Nevertheless, a unified paradigm has emerged for training LLMs: pre-train on an enormous corpus of diverse documents, ranging from 250B Biderman et al. (2023) to 15T AI@Meta (2024) tokens, followed by an alignment stage to make the model more useful and performative for various tasks.
Figure 1: Illustration of the experimental scheme.
@@
Based on this paradigm, work has focused on improving these two stages. Work to improve pre-trained models includes larger training sets Hoffmann et al. (2022); AI@Meta (2024); Touvron et al. (2023), different data selection mechanisms Xia et al. (2024), higher quality data Zhou et al. (2024), and various model architectures Su et al. (2024); Touvron et al. (2023). Meanwhile, research on model alignment includes different training objectives Rafailov et al. (2024); Schulman et al. (2017), new datasets Narayanan and Aepli (2024), more efficient training Hu et al. (2021); Dettmers et al. (2024) and safety tuning Bianchi et al. (2023). The alignment stage usually involves either supervised fine-tuning for specific tasks or instruction fine-tuning for general-purpose usage. Regardless, fine-tuning (almost always) comes at the end of pre-training and yields remarkable improvements on downstream tasks Touvron et al. (2023); Groeneveld et al. (2024). Consequently, the benefits of each stage are largely explored independently, with improvements to pretraining being orthogonal to benefits from model alignment.
@@
Rather than explore these two training regimes independently, we question: how do model pretraining and fine-tuning interact to affect the resulting model? Does more pre-training hinder better fine-tuning results? What does the model learn and forget during pre-training as well as fine-tuning? Answering these questions requires us to examine how models learn during pre-training and how this affects fine-tuning. Therefore, we fine-tune multiple pre-training checkpoints of a large language model (Figure 1), evaluating each checkpoint and its fine-tuned version on downstream evaluation sets. We track model abilities during pre-training and compare them to improvements achievable after fine-tuning at the corresponding pre-training step. We explore both supervised and instruction fine-tuning, testing the models’ memorization and forgetting when learning specific tasks and serving as general-purpose language-AI tools. To the best of our knowledge, we are the first to explore fine-tuning intermediate model checkpoints.
@@
Our experiments yield insights into LLM training. We find that (1) continued pre-training can improve a model in ways that are only revealed after fine-tuning (§5); (2) tasks for which the model already performs well during pre-training benefit much less from fine-tuning than those where the model does not demonstrate capabilities (§4, §5); (3) although supervised fine-tuning can improve performance on in-distribution tasks, it can also cause the model to forget domain knowledge or tasks that it was previously capable of solving (§6); (4) fine-tuned models show high sensitivity to evaluation prompts, but this sensitivity can be alleviated by more pre-training (§6). Our findings provide insights into model training and can inform methods for both pre-training and fine-tuning. Furthermore, our work shows the value of analyzing the training dynamics, in addition to analyzing the final LLM, as an aspect of interpretability, and we encourage model developers to release these checkpoints to aid future studies.
@@ 2 Background: Model Training
We begin with a brief survey of the core components of LLM training: pre-training, fine-tuning, and instruction fine-tuning. We also discuss the related topic of in-context learning as well as different efficient fine-tuning strategies.
@@
We use “model alignment” as a general term for techniques that align a model with a desired behavior, which can be accomplished by fine-tuning models after pretraining. The term is also associated with other definitions Shen et al. (2024). We also note several related studies that explore training dynamics to understand model behavior Tirumala et al. (2022); Chen et al. (2023); Tian et al. (2023). With this in mind, we conduct an empirical study on how the amount of pre-training affects the effectiveness of fine-tuning.
Pre-training
@@
The first step of training a LLM is pre-training on a massive text corpus Achiam et al. (2023); Touvron et al. (2023); Groeneveld et al. (2024). For decoder-only models in the GPT family, the subject of our paper, work since the introduction of GPT-2 Radford et al. (2019) has focused on scaling up model training. Initial work increased model size to hundreds of billions of parameters Brown et al. (2020); Rae et al. (2021); Chowdhery et al. (2023), along with explorations in model size, training corpus size, and training data characteristics Hoffmann et al. (2022); Gururangan et al. (2020). Since the push towards large models, work has shifted to increasing the amount of pre-training data Computer (2023); Soldaini et al. (2024), with new models now reaching 15 trillion tokens AI@Meta (2024). Studies of model performance on various tasks at different model sizes introduced the idea of emergent model abilities Wei et al. (2022), with new model abilities being revealed as model training grows.
We also recognize a particularly important trend for this paper: model openness. Early LLMs were proprietary models accessible only through an API. The first large open model, Bloom Bloom Ström et al. (2023), allowed widespread LLM evaluation. Subsequent open models, such as OPT Zhang et al. (2022), LLaMA Touvron et al. (2023); Keles and Bayraklı (2024) and others Biderman et al. (2023); Gururangan et al. (2023); Almazrouei et al. (2023), have become the norm. In this paper, we study OLMo Groeneveld et al. (2024), one of the only models to release individual pre-training checkpoints.
Fine-Tuning
@@
Early work on instruction fine-tuning using reinforcement learning with human feedback (RLHF) Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022) demonstrates the dramatic effect that model alignment could have on a pre-training model. When a specific task of interest has been identified, supervised fine-tuning can improve a pre-trained model. Task-agnostic tuning became popularized with the advent of T5 models (Raffel et al., 2020), where a pre-trained LLM is tuned using a general text-to-text solution. When multiple tasks are given to the model, the model is commonly given a task-specific prefix or an instruction along with the task input, leading to the development of various methods of prefix tuning Li and Liang (2021) and instruction tuning Wei et al. (2021); Mishra et al. (2022); Victor et al. (2022).
Instruction Fine-Tuning
Instruction fine-tuning is preferred when more general model behaviors are desired. Popularized through reinforcement-learning with human feedback (RLHF) Christiano et al. (2017); Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022) and reinforcement-learning with AI feedback (RLAIF) Lee et al. (2023), these methods utilize a reward model to simulate human feedback. Others explore human preference tuning without a reward model Rafailov et al. (2024); Song et al. (2024); Xu et al. (2024), or study the effects of these tuning methods (Shen et al., 2024; Perez et al., 2023). Sharma et al. (2024) show that supervised fine-tuning can lead to similar performance as RLAIF.
In-Context Learning
While not the subject of this paper since it does not make changes to model parameters, in-context learning utilizes a small amount of supervised data to improve model performance. ICL, also called few-shot learning, is also used as an evaluation strategy where the model is given a prompt composed of examples of tasks expected to be solved. The underlying model is evaluated based on its response to the input. ICL can benefit from a larger context window that adds more examples, which can spur work on the development of model quantization techniques Dettmers et al. (2022) and the alleviation of hardware constraints Brown et al. (2020); Xie et al. (2021); Min et al. (2022).
Fine-Tuning Techniques
While model pre-training can be done by a few groups with large resources interested in developing new models, fine-tuning depends on the task and is of broad interest. Therefore, many techniques facilitate time-, memory-, and data-efficient model training through parameter-efficient fine-tuning (PEFT) Hu et al. (2021); Liu et al. (2021, 2023), quantization Jacob et al. (2018); Dettmers et al. (2022, 2024), and specialized data filtering Xia et al. (2024); Zhou et al. (2024); Attendu and Corbeil (2023). This paper focuses specifically on full-parameter fine-tuning, while our findings suggest the potential for data-efficient and budget-friendly training by understanding the critical turning point of model training. Our findings are closely related to the recent study on phase transition of model training Olsson et al. (2022); Wei et al. (2022); Chen et al. (2023).
@@ 3 Experimental Setup
In this section, we describe the model and datasets used. The hyperparameter tuning procedure and setup for each fine-tuning setting can be found in Appendix A.
3.1 Model Choice
Our paper considers OLMo-1B Groeneveld et al. (2024), a high-performing open-source large language model. Ideally, we would evaluate multiple models, but OLMo is the only model to release intermediate pre-training checkpoints, and thus the only model that supports our analysis [^2] https://github.com/allenai/OLMo/tree/main/checkpoints  [^3] We also experimented with RedPajama-INCITE (https://www.together.ai/blog/redpajama-models-v1), which is the only other model to release checkpoints. After extensive experiments, we found it performed worse than OLMo, given the training data available, and did not support our analysis. Several other models claim to release training checkpoints but have not done so.. Despite being the only open model with training checkpoints, it fortunately has several desirable properties. First, the model is fully open, including the training details, pre-training data, and fine-tuning data. Second, the smaller model size allows us to train a model efficiently on a single A100 GPU. While evaluating a larger model would be desirable, we limit our study to the 1B model given the much larger computational demand of multi-GPU training. Our detailed analysis required significant GPU resources, which would have been prohibitive with a larger model. We also note that OLMo-1B compares very favorably to the larger version, and recent work has shown that small models can compete with larger ones Riviere et al. (2024).
We select model pre-training checkpoints uniformly from the pre-training history along with the first and the final checkpoints.
3.2 Training Procedure
We fine-tune each of the selected model checkpoints using two different procedures to create fine-tuned models: supervised fine-tuning and instruction tuning. The supervised fine-tuning is conducted separately for each model checkpoint and dataset, while the instructing fine-tuning is done once using the instruction dataset. The instruction-tuned model is evaluated on a suite of LLM benchmarks.
Supervised Fine-tuning
We adapt the dataset choice from Yang et al., 2024 for supervised fine-tuning. For each in-domain dataset, one to two cross-domain evaluation datasets are supplied. Each pre-training checkpoint is fully fine-tuned for 3 epochs with a batch size of 8 and learning rates resulting from minimal hyperparameter tuning. Each task is formatted using a default prompt-completion format (Table 4).
Instruction Fine-Tuning
We instruction-tune the model on TÜLU Ivison et al. (2023), following the decision of Groeneveld et al., 2024. Each model checkpoint is fully fine-tuned for 5 epochs with a batch size of 8 and a learning rate of $2\times 10^{-6}$.
@@ 3.2 Training Procedure
Table 1: Dataset information. For Generation tasks, ROUGE-L is used as evaluation metric, and accuracy is used for classification tasks.
@@ 3.3 Evaluation
The evaluation challenge is to select a representative number of datasets for different types of tasks to test model abilities, recognizing that each dataset requires evaluating each model checkpoint and its fine-tuned counterparts. We also select datasets based on the availability of in-domain and out-of-domain samples.
Datasets
@@
The datasets are summarized in Table  1. We evaluate the model with an in-domain test set and one or two out-of-domain test sets for each of the supervised fine-tuning tasks. We conduct experiments on the tasks of summary generation Narayan et al. (2018); Hasan et al. (2021); Hermann et al. (2015), question generation Sap et al. (2019); Xiong et al. (2019); Welbl et al. (2017), natural language inference Williams et al. (2018); Wang et al. (2018); Dagan et al. (2006); Bar Haim et al. (2006); Giampiccolo et al. (2007); Bentivogli et al. (2009), and paraphrase detection Zhang et al. (2019); Wang et al. (2018); Agirre et al. (2007). Each training set is sub-sampled to a size of 6,000 for fair comparisons.
@@
In instruction fine-tuning, we base our downstream evaluation settings on Groeneveld et al., 2024, as OLMo is found to have stable performance on these datasets. The instruction-tuned models are evaluated on ARC (both arc easy and arc challenge) Clark et al. (2018), OpenbookQA Mihaylov et al. (2018), Hellaswag Zellers et al. (2019), BoolQ Clark et al. (2019), and SciQ Welbl et al. (2017).
Metrics
@@
We use accuracy Pedregosa et al. (2011) for classification tasks and Rouge-L Lin (2004) for generation tasks. We set the maximum amount of newly generated tokens to 5 for classification tasks and 60 for generation tasks. Outputs are generated with greedy decoding. For classification tasks, we experiment with both constrained decoding and logit-based predictions. We find the best performance by selecting the label with the highest logit of its first subtoken (Appendix B).

### v5

@@ Abstract
Large language model development relies on the pre-train-then-align paradigm, in which the model is typically pre-trained on a large text corpus and undergoes a tuning stage to align the model with human preference or downstream tasks. We investigate the relationship between pre-training and supervised fine-tuning by considering multiple tasks as well as different pre-trained model checkpoints. Our results on 18 datasets and two models suggest that i) although the model benefits significantly through supervised fine-tuning, it may forget previously known domain knowledge and tasks that are not seen during fine-tuning; ii) the model exhibits high sensitivity to evaluation prompts after supervised fine-tuning, but this sensitivity can be alleviated through further pre-training; iii) continual pre-training improves the model in a latent way that manifests after fi ne-tuning; iv) The model can already solve some tasks after pre-training, while fine-tuning most benefits datasets where the model does not show capability during pre-training.  [^1] Code, results, and data to reproduce the experiments are available at github.com/KaiserWhoLearns/AmuroCharPTFTRelationship. All the model checkpoints resulting from this work are available at huggingface.co/KaiserWhoLearns/PTvsSFT_OLMo1b
@@ 1 Introduction
The rise of large language models (LLMs) as a general-purpose tool for a diverse range of natural language processing tasks has dramatically transformed the field, introducing new paradigms for data collection and model training (Brown et al., 2020, Biderman et al., 2023, Touvron et al., 2023, Jiang et al., 2023, Chowdhery et al., 2023, Groeneveld et al., 2024, Wang et al., 2024, inter alia). Numerous models, training methods, datasets, and evaluation methods continue to be developed on an ongoing basis. Nevertheless, a unified paradigm has emerged for training LLMs: pre-train on an enormous corpus of diverse documents, ranging from 250B (Biderman et al., 2023) to 15T (AI@Meta, 2024) tokens, followed by an alignment stage to make the model more useful and performative for various tasks.
@@
Based on this paradigm, work has focused on improving these two stages. Work to improve pre-trained models includes larger training sets (Hoffmann et al., 2022; AI@Meta, 2024; Touvron et al., 2023), different data selection mechanisms (Xia et al., 2024), higher quality data (Zhou et al., 2024), and various model architectures (Su et al., 2024; Touvron et al., 2023). Meanwhile, research on model alignment includes different training objectives (Rafailov et al., 2024; Schulman et al., 2017), new datasets (Narayanan and Aepli, 2024), more efficient training (Hu et al., 2021; Dettmers et al., 2024) and safety tuning (Bianchi et al., 2023). The alignment stage usually involves either supervised fine-tuning for specific tasks or instruction fine-tuning for general-purpose usage. Regardless, fine-tuning (almost always) comes at the end of pre-training and yields remarkable improvements on downstream tasks (Touvron et al., 2023; Groeneveld et al., 2024).
Figure 1: Illustration of the experimental scheme. Intermediate pre-training checkpoints are fine-tuned on different datasets.
Consequently, the benefits of each stage are largely explored independently, with improvements to pretraining being orthogonal to benefits from model alignment.
@@
Rather than exploring these two training regimes independently, we ask: What does the model learn and forget during pre-training and fine-tuning? Specifically, how do pretraining and fine-tuning interact to produce the resulting model? Does more pre-training hinder better fine-tuning results? Answering these questions requires us to examine how models learn during pre-training and how this affects fine-tuning. Therefore, we begin by fine-tuning two language models under a variety of conditions to determine how fine-tuning affects model behavior. We explore both supervised and instruction fine-tuning, testing the models’ memorization and forgetting when learning specific tasks and serving as general-purpose language-AI tools. We then explore the affect of pre-training on these behaviors by fine-tuning multiple pre-training checkpoints of a large language model (Figure 1), evaluating each checkpoint and its fine-tuned variant on downstream evaluation sets. We track model abilities during pre-training and compare them to improvements achieved after fine-tuning at the corresponding pre-training step. [^2] While we believe that we were the first to explore these issues through intermediate model checkpoints, recently released work has also utilized pre-training checkpoints and are highlighted in Section 8.
@@
Our experiments yield the following insights into LLM training: (1) although supervised fine-tuning can improve performance on in-distribution tasks, it can also cause the model to forget domain knowledge or tasks that it was previously capable of solving (§4); (2) fine-tuned models show high sensitivity to evaluation prompts, but this sensitivity can be alleviated by more pre-training (§4); (3) continued pre-training can improve a model in ways that are only revealed after fine-tuning (§6); (4) tasks for which the model already performs well during pre-training benefit much less from fine-tuning than those where the model does not demonstrate capabilities (§5, §6);
Our findings provide insights into model training and can inform methods for both pre-training and fine-tuning. Furthermore, our work shows the value of analyzing the training dynamics, in addition to analyzing the final checkpoint of an LLM, as an aspect of interpretability, and we encourage model developers to release these checkpoints to aid future studies.
@@ 2 Background: Model Training
We use “model alignment” as a general term for techniques that align a model with a desired behavior, often accomplished by fine-tuning models after pretraining. The term is also associated with other definitions (Shen et al., 2024).
@@
We begin with a brief survey of the core components of LLM training: pre-training, fine-tuning, and instruction fine-tuning. The first step of training an LLM is pre-training on a massive text corpus (Achiam et al., 2023; Touvron et al., 2023; Groeneveld et al., 2024). Initial work increased model size to hundreds of billions of parameters (Brown et al., 2020; Rae et al., 2021; Chowdhery et al., 2023), along with explorations in model size, training corpus size, and training data characteristics (Radford et al., 2019; Hoffmann et al., 2022; Gururangan et al., 2020). Other work increased the amount of pre-training data (Computer, 2023; Soldaini et al., 2024), with new models now reaching 15 trillion tokens (AI@Meta, 2024).
@@
After the pre-training stage, when a specific task of interest has been identified, supervised fine-tuning can improve a pre-trained model. Task-agnostic tuning became popularized with the advent of T5 models (Raffel et al., 2020), where a pre-trained LLM is tuned using a general text-to-text solution. Instruction fine-tuning is preferred when more general model behaviors are desired. When multiple tasks are given to the model, the model is commonly given a task-specific prefix or an instruction along with the task input, leading to the development of various methods of prefix tuning (Li and Liang, 2021) and instruction tuning (Wei et al., 2021; Mishra et al., 2022; Victor et al., 2022).
Other works explore human preference tuning with or without a reward model (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Rafailov et al., 2024; Song et al., 2024; Xu et al., 2024). In-context learning utilizes a small amount of supervised data to improve model performance without updating the parameters. In this work, we focus specifically on single-task supervised fine-tuning and multi-task instruction tuning.
@@ 3 Experimental Setup
In this section, we describe the models and datasets used. The hyperparameter tuning procedure and setup for each fine-tuning setting can be found in Appendix B.
@@
Table 1: Dataset information. For Generation tasks, ROUGE-L is used as evaluation metric, and accuracy is used for classification tasks. ID = In-domain, OOD = Out-of-domain.
3.1 Model Choice
We consider two open models of different architectures and scales: Llama3-8B (AI@Meta, 2024) and OLMo-1B (Groeneveld et al., 2024). To minimize potential confounding factors such as multilingual ability and double descent (Belkin et al., 2019; Caballero et al., 2022; Schaeffer et al., 2023), we exclusively select models predominantly pre-trained in English and incorporate significantly more pre-trained tokens than the number of parameters. We do not include models trained in a multi-stage manner to ensure uniformity of the tokens seen by the model during pre-training. Some of our experiments consider intermediate pre-training checkpoints. We select checkpoints uniformly by the number of tokens seen from the pre-training history along with the first and the final checkpoints. Unfortunately, very few large language models release intermediate pre-training checkpoints (summarized in Table 2). Further consideration and reasoning of model selection are included in Appendix A.
3.2 Training Procedure
We fully fine-tune each of the selected model checkpoints using two different procedures to create fine-tuned models: supervised fine-tuning and instruction tuning. The supervised fine-tuning is conducted separately for each model checkpoint and dataset, while the instruction fine-tuning is done once using the instruction dataset. The instruction-tuned model is evaluated on a suite of LLM benchmarks. All experiments are conducted on two Nvidia 80GB A100, with a total cost of approximately 1100 GPU hours. The detailed number of GPU hours consumed for each experiment is included in Appendix E.
Supervised Fine-tuning. We adapt the datasets from Yang et al. (2024) for supervised fine-tuning. For each in-domain dataset, one to two cross-domain evaluation datasets are supplied. OLMo-1B is fully fine-tuned for 3 epochs with a batch size of 8, while Llama3-8B is fine-tuned with a batch size of 16 and 2 training epochs. Both models are trained with learning rates resulting from minimal hyperparameter tuning (Appendix B). Each task is formatted using a default prompt-completion format (Table 5).
Instruction Fine-Tuning. We instruction-tune the model on TÜLU (Ivison et al., 2023), following the decision of Groeneveld et al., 2024. Each model checkpoint is fully fine-tuned for 5 epochs with a batch size of 8 and a learning rate of $2\times 10^{-6}$.
@@ 3.3 Evaluation
For each model, we conduct a few-shot evaluation with a shot size of 4, after examining with shot size in {0, 2, 4, 6}.
((a)) MNLI matched
((b)) SocialIQa
Figure 3: LLAMA3-8B performance with different task format. Instruct and Default always lead to highest evaluation results.
@@
Datasets. The datasets are summarized in Table  1 and data licenses are in Table 7. We evaluate the model with an in-domain test set and one or two out-of-domain test sets for each of the supervised fine-tuning tasks. We conduct experiments on the tasks of summary generation (Narayan et al., 2018; Hasan et al., 2021; Hermann et al., 2015), question generation (Sap et al., 2019; Xiong et al., 2019; Welbl et al., 2017), natural language inference (Williams et al., 2018; Wang et al., 2018; Dagan et al., 2006; Bar Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009), and paraphrase detection (Zhang et al., 2019; Wang et al., 2018; Agirre et al., 2007). We subsample 6,000 training instances for each set to ensure a fair comparison.
@@
In instruction fine-tuning, we base our downstream evaluation settings on Groeneveld et al. (2024), as OLMo is found to have stable performance on these datasets. The instruction-tuned models are evaluated on ARC (both arc easy and arc challenge) (Clark et al., 2018), OpenbookQA (Mihaylov et al., 2018), Hellaswag (Zellers et al., 2019), BoolQ (Clark et al., 2019), and SciQ (Welbl et al., 2017).
@@
Metrics. We use accuracy (Pedregosa et al., 2011) for classification tasks and Rouge-L (Lin, 2004) for generation tasks. The maximum amount of newly generated tokens is set to 5 for classification tasks and 60 for generation tasks. Outputs are generated with greedy decoding. For classification tasks, we experiment with both constrained decoding and logit-based predictions. We find the best performance by selecting the label with the highest logit of its first subtoken (Appendix C).
4 Supervised Fine-Tuning: What does the model learn and forget?
We begin our analysis with the supervised fine-tuning process to understand the downstream results of the training process. Specifically, we explore three dimensions: task format, task transfer, and domain knowledge. In each case, we fine-tune both final checkpoints and intermediate pre-training checkpoints to understand the relationship between pre-training and fine-tuning.
4.1 Task Format
LLMs can be extremely sensitive to prompt perturbation in few-shot settings (Sclar et al., 2023; Leidinger et al., 2023; Salinas and Morstatter, 2024; Wahle et al., 2024). We hypothesize that fine-tuning fits the model to a specific task format, resulting in higher performance when the evaluation set matches this format. To test this hypothesis, we vary the task format to either match the training format, use a different format, or rely on instructions.
We carefully construct three different prompt formats for the following settings. 1) Default is the same format used for training, where we expect the model to benefit from learning the task format; 2) IO format, by contrast, reflects a common way of performing supervised fine-tuning by incorporating only unprocessed input and output; 3) Instruct uses a human-readable instruction template to format the input. Table 5 in the Appendix shows an example of each format. The performance of Llama3-8B with different task formats is shown in Figure 3. Checkpoint performance on OLMo before and after fine-tuning is shown in Figure 2.
Across both models, IO format leads to the least favorable performance, as the only task-specific information in this format is included in the evaluation shots. Model reports similar performance when evaluated with the default and instruct format, aligning with the findings in Hewitt et al. (2024) that the models retain their instruction-following ability after fine-tuning without instructions. However, in the early pre-training steps, aligning the task format with fine-tuning data plays a crucial role (Figure 2), suggesting that the instruction-following ability has not yet been developed. In this view, fine-tuning teaches the model how to format a response for the task, while further pretraining enhances the instruction-following ability. In other words, the instruction provides a directed prior for the model to behave in a certain way.
((a)) Paws $\rightarrow$ QQP
((b)) MNLI $\rightarrow$ GPT3NLI
Figure 5: Ratio of out-of-domain performance change for each task, averaged across checkpoints.
4.2 Domain Knowledge
We next explore how the domain-generalization ability is affected by fine-tuning by inspecting whether the model forgets the domain knowledge after fine-tuning on a different domain. An example of OOD model performance is shown in Figure 4, and the mean ratio of change by datasets is presented in Figure 5 and Figure 15.
The models do not benefit equally from the in-domain fine-tuning: Llama shows subtle benefits on question generation tasks, while not benefiting at all on the other tasks (Figure 15). Across OLMo training history (Figure 5), NLI datasets experience a boost when fine-tuning on MNLI, while fine-tuning on Paws is detrimental to other paraphrase detection datasets. This suggests that both forgetting and learning are happening in fine-tuning: the model learns to perform the task with in-domain knowledge, but it may, in turn, forget information more distant from what is learned in fine-tuning. Furthermore, under the same task, the amount of general-purpose pre-training may not affect the model’s reaction to out-of-domain knowledge. Questions remain, however, about whether domain-specific continual pre-training or continual pretraining on similarly distributed data would bring different conclusions, which requires further study of pre-training dynamics.